# Partially Retargeted Balancing Weights for Causal Effect Estimation Under Positivity Violations

### v1

@@ Introduction
We can alternatively explore term (8a) for our proposed estimator from a design-based, rather than modeling, prospective; an alternative simplification of term (8a) is $(\bm{\lambda}_{1}^{T}-\bm{\lambda}_{0}^{T})\left\{\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]g(\bm{X}_{i})\right$. $\left.-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\}$. From a design-based perspective, this term tends to be small when $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ is small. However, in Problem (5) we explicitly do not constrain $\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})=0$ in order to yield a solution when there is a lack of overlap. Thus, $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w^{*}(\bm{X}_{i})I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ will be non-zero when weights are derived using Problem (5). Despite this, the error of our proposed estimator with respect to the ATE is due to differences between the weighted and sample populations of $g(\bm{x})$, rather than the imbalance between the weighted treated and control populations. Thus, when $c(\bm{x})$ does not contain all treatment effect modifiers, the estimator has minimal error with respect to the average treatment effect estimand over a modified population of $g(\bm{x})$, which can be characterized by $\frac{1}{n}\sum_{i=1}^{n}w_{i}Z_{i}g(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})g(\bm{X}_{i})$ in comparison to the overall sample mean. While the ATO population can also be characterized through weighted sample means, it can be challenging to conceptualize this modified population especially with high-dimensional covariates. In contrast to both overlap weights and our proposed estimator, the minimal weights estimator does not target another estimand of interest and its bias with respect to the ATE is predominantly due to covariate imbalance. We provide further guidance and discussion on how to select $g(\bm{x})$ given these model- and design-based perspectives in Section 6 and supplementary Section 4.1.
Asymptotic properties of the proposed estimator
We derive asymptotic results of the solution to Equation (6) which we then use to obtain asymptotic results of our proposed estimator, $\tau_{w^{*}}$, under various conditions. All proofs of results in this section are in supplementary Section 1. We first require an additional regularity condition, similar to those in Hirano et al., (2003); Chan et al., (2016) and Källberg and Waernbaum, (2023), on the set balancing functions, $b(\bm{x})$.
Assumption 3: Balance function regularity conditions.
For the remainder of this section, we notate $w_{i}^{*}$ as $w^{*}(\bm{X}_{i})$ to explicitly show the relationship to $\bm{X}_{i}$. Given Assumptions Assumption 1: Strong ignorability, Assumption 2: Relaxed positivity, and Assumption 3: Balance function regularity conditions, we show that $w_{1}^{*}(\bm{X}_{i})$ and $w_{0}^{*}(\bm{X}_{i})$ are consistent for their asymptotic counters, $\tilde{w}_{1}(\bm{X}_{i})$ and $\tilde{w}_{0}(\bm{X}_{i})$ that uniquely solve the population version of Equation 6 (supplementary Section 1, Proposition 5). Since Problem (5) implies a single propensity score model, our proposed estimator is doubly robust as stated below in Theorem 3; in contrast, $\hat{\tau}_{w^{bw}}$ is not doubly robust (Chan et al.,, 2016).
Theorem 3.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 3: Balance function regularity conditions hold. Assume that the expectation of $c(\bm{x})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then our proposed estimator, $\hat{\tau}_{w^{*}}$, is doubly robust in the sense that if either 1) Assumption Assumption 2: Positivity holds and $e(\bm{x})$ satisfies conditions (7a)-(7c) or; 2) Assumption Assumption 2: Relaxed positivity holds, $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$ then $\hat{\tau}_{w^{*}}$ is consistent for the ATE.
While $\hat{\tau}_{w}^{*}$ is doubly robust, we note that $e(\bm{x})$ will rarely satisfy the restrictive conditions (7a)-(7c). Still, our proposed estimator is consistent under the relaxed positivity assumption if $c(\bm{x})$ contains all effect modifiers. When $c(\bm{x})$ does not contain all effect modifiers, our proposed estimator is consistent for the modified estimand $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]-E[E[Y|Z=0,c(\bm{X}),\tilde{w}_{0}(\bm{X})\{1-e(\bm{X})\}g(\bm{X})]]$ as stated in Corollary 3.1.
Corollary 3.1.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 2: Relaxed positivity hold. Assume that the expectation of $c(\bm{x})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then our proposed estimator, $\hat{\tau}_{w^{*}}$ is consistent for $\tau_{g,\tilde{w}}$ if $e(\bm{x})$ does not satisfy conditions (7a)-(7c) and $\mu_{z}(\bm{x})$, $\tau(\bm{x})$ are linear in $c(\bm{x})$ and $g(\bm{x})$.
Note that $\tau_{g,\tilde{w}}$ is indeed a causal estimand because $E[\tilde{w}_{1}(\bm{X}_{i})e(\bm{X}_{i})g(\bm{X_{i}})]=E[\tilde{w}(\bm{X}_{i})_{0}\{1-e(\bm{X}_{i})\}g(\bm{X_{i}})]$ by $\tilde{w}$ satisfying the first order conditions of the population version of Equation (6). Thus, we can also write $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]-E[E[Y_{i}|Z_{i}=0,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]$.
When positivity does hold and the implied propensity score and outcome models of Problem 5 are correctly specified, our proposed estimator is semiparametrically efficient as stated in Theorem 4.
Theorem 4.
Suppose Assumptions 11, 22, and 33 hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ satisfies conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt})$
where $V_{opt}=E\left[\frac{\text{Var}(Y(1)|\bm{X})}{e(\bm{X})}+\frac{\text{Var}(Y(0)|\bm{X})}{1-e(\bm{X})}+\{\tau(\bm{X})-\tau\}^{2}\right]$ is the semiparametric efficiency bound (Hahn,, 1998).
However, it is rare that the implied propensity score model is correctly specified and our proposed estimator is motivated by scenarios where Assumption Assumption 2: Positivity does not hold. We thus derive and present the asymptotic variance of our estimator when only the outcome models are correctly specified in Corollary 4.1.
Corollary 4.1.
Suppose Assumptions 11, $2^{*}$, and 33 hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ does not satisfy conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then,
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt}^{*})$
where $V_{opt}^{*}=E[\tilde{w}_{1}(\bm{X})^{2}e(\bm{X})\text{Var}\{Y(1)|\bm{X}_{i}\}+\tilde{w}_{0}(\bm{X})^{2}\{1-e(\bm{X})\}\text{Var}\{Y(0)|\bm{X}_{i}\}+\{\tau(\bm{X})-\tau\}^{2}]$
Of key importance, this variance result only requires the relaxed positivity assumption, rather than the standard positivity assumption. However, to compare $V_{opt}$ and $V^{*}_{opt}$, consider the case where positivity does hold. Since the asymptotic version of the weights tend to be less extreme (i.e., $\tilde{w}_{1}(\bm{X}_{i})\geq e(\bm{X}_{i})$ and $\tilde{w}_{0}(\bm{X}_{i})\leq 1-e(\bm{X}_{i})$) than IPW through the restriction on the implied propensity score models, generally $V^{*}_{opt}\leq V_{opt}$. Thus, our proposed estimator does indeed have reduced asymptotic variance, especially in scenarios where $e(\bm{x})$ is close to zero or positivity does not hold. While intuitively as the $g(\bm{x})$ set gets larger, estimator variance decreases, Corollary 4.1 formalizes this intuition; as the $g(\bm{x})$ set gets larger, $\tilde{w}_{z}(\bm{x})$ become less extreme and correspondingly $V^{*}_{opt}$ becomes smaller.
@@
Our proposed constrained optimization problem for deriving weights for an ATE estimator can be readily extended to variety of additional estimation problems. We discuss extensions to the ATT and transporting effects below. Discussion of causal effect estimation with more than two treatment groups, weighted average treatment effects, and distributional balancing weights are in supplementary Section 3.
@@
When deriving either IPW or direct balancing weights for the ATT, only weights for the control group are derived such that the corresponding estimator is $\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})Y_{i}$. However, when adapting Problem (5) for ATT estimation, weights must be derived for both the treated and control group. To derive these weights, $\{w_{i}^{*ATT}\}$, we can simply replace the right-hand side of constraint (5b) with $\frac{1}{n}\sum_{i=1}^{n}Z_{i}c_{k}(\bm{X}_{i})$ and compute Equation (4) with these weights. The design-based and model-based interpretations of the sets $c(\bm{x})$ and $g(\bm{x})$ remain the same as described in Section 3.2. However, the covariates that may be effect modifiers for the entire sample population potentially could not be for the treated population. Further, the methods for a data-driven selection of $g(\bm{x})$ proposed in Section 6 can be used here as well.
@@
While clinical trials yield ‘gold standard’ causal effect estimates through randomization, the inclusion/exclusion trial criteria is often restrictive such that the trial population may not reflect a population of interest. In these cases, it is desirable to transport the effect to a more meaningful population. However, there is often a lack of overlap between the restrictive trial population and the target population such that methods similar to propensity score trimming or overlap weights must be used. Our procedure allows the analyst to improve overlap by modifying the target population for only a subset of covariates, rather than modifying all covariates or discarding observations, which is inconsistent with the goal of estimating the effect for a meaningful population. Before describing the adaptation of Problem 5 for estimating a transported effect, we introduce some additional notation. Let $R_{i}=r$ be an indicator of whether an observation belongs to the trial population ($r=1$) or the target population ($r=0$). We only observe $Y_{i}$ and $Z_{i}$ when $R_{i}=1$ but observe $\bm{X}_{i}$ for both populations. Let $n_{r}$ indicate the sample size in population $r$. Given additional assumptions outlined in Dahabreh et al., (2020), the transported effect estimand is identifiable with observed data. We propose the following modification of Problem (5) for transportability,
@@
Then, our proposed estimator for the transported effect is $\tau_{w^{*t}}=\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}Z_{i}R_{i}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}(1-Z_{i})R_{i}Y_{i}$. The interpretation of $c(\bm{x})$ and $g(\bm{x})$ remains the same as described in Section 3.2 however the selection of $g(\bm{x})$ may be of even more importance as maintaining as much of the target population as possible is of key interest when transporting effects.
@@
To select the optimal $g(\bm{x})$ set, in terms of error with respect to the ATE, one approach is to examine all possible combinations of functions of covariates in the set $g(\bm{x})$ and select the set that 1) yields a solution to Problem (5); and 2) minimizes term (8a). However, examining all possible covariate subsets is computationally prohibitive, especially as covariate dimension increases. Therefore, we propose a greedy algorithm (Algorithm 1) for identifying an approximately optimal $g(\bm{x})$ subset according to a specific metric of interest aligned with either the model- or design- based perspective on estimator error. We also propose an additional algorithm (supplementary Algorithm 1) that is less computationally intensive than Algorithm 1; we found these algorithms resulted in estimators with similar performance when there was no correlation between covariates (supplementary Figure 2). However, we generally recommend the use of Algorithm 1 if it is not computationally prohibitive.
@@
1:Inputs: Set of functions of covariates $b_{1}(\bm{X}),\ldots,b_{J}(\bm{X})$ and a metric $m$ that can be computed for each $b_{j}(\bm{X})$
@@
3:while $w^{*}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
4:  Compute metric $m$ for $\{b_{1}(\bm{X}),\ldots,b_{j}(\bm{X})\}\setminus b(\bm{X})^{-}\}$ and identify $b^{*}(\bm{X})$, the covariate with the minimum or maximum value of the metric $m$
@@
6:  Solve for $w^{*}$ in Problem (5) for $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{1}(\bm{X}),\ldots,b_{j}\}\setminus b(\bm{X})^{-}$

### v2

@@ Introduction
We can also explore term (8a) for our proposed estimator from a design-based, rather than modeling, prospective; an alternative simplification of term (8a) is $(\bm{\lambda}_{1}^{T}-\bm{\lambda}_{0}^{T})\left\{\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]\right$. $\left.\times g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\}$. From a design-based perspective, this term tends to be small when $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ is small. However, in Problem (5) we explicitly do not constrain $\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})=0$ in order to yield a solution when there is a lack of overlap. Thus, $1^{T}\left\lvert\frac{1}{n}\sum_{i=1}^{n}w_{i}I[Z_{i}=z]g(\bm{X}_{i})-\frac{1}{n}\sum_{i=1}^{n}g(\bm{X}_{i})\right\rvert$ will be non-zero when weights are derived using Problem (5). Despite this, the error of our proposed estimator with respect to the ATE is due to differences between the weighted and sample population of $g(\bm{x})$, rather than the imbalance between the weighted treated and control populations. Thus, when $c(\bm{x})$ does not contain all treatment effect modifiers, our proposed estimator has minimal error with respect to the average treatment effect estimand over a modified population of $g(\bm{x})$, which can be characterized by $\frac{1}{n}\sum_{i=1}^{n}w^{*}_{i}Z_{i}g(\bm{X}_{i})=\frac{1}{n}\sum_{i=1}^{n}w^{*}_{i}(1-Z_{i})g(\bm{X}_{i})$. While the ATO population can also be characterized through weighted sample means, it can be challenging to conceptualize this modified population especially with high-dimensional covariates. We provide further guidance and discussion on how to select $g(\bm{x})$ given these model- and design-based perspectives in Section 5 and supplementary Section 5.1.
@@
Our proposed constrained optimization problem for deriving weights for ATE estimation can be readily extended to variety of additional estimation problems. We discuss extensions to the ATT and transporting effects below. Extensions to effect estimation with more than two treatment groups, weighted average treatment effects, and distributional balancing weights are discussed in supplementary Section 3.
@@
When deriving either IPW or direct balancing weights for the ATT, only weights for the control group are derived such that the corresponding estimator is $\frac{1}{n}\sum_{i=1}^{n}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w_{i}(1-Z_{i})Y_{i}$. However, when adapting Problem (5) for ATT estimation, weights must be derived for both the treated and control group. To derive these weights, $\{w_{i}^{*ATT}\}$, we can simply replace the right-hand side of constraint (5b) with $\frac{1}{n}\sum_{i=1}^{n}Z_{i}c_{k}(\bm{X}_{i})$ and compute Equation (4) with these weights. The design-based and model-based interpretations of the sets $c(\bm{x})$ and $g(\bm{x})$ remain the same as described in Section 3.2. However, the covariates that are effect modifiers may differ between the entire sample population and the treated population.
@@
While clinical trials yield “gold standard” causal effect estimates through randomization, trial inclusion/exclusion criteria is often restrictive such that the trial population may not reflect a clinically relevant population. In these cases, it is desirable to transport the effect to a more meaningful population. However, there is often a lack of overlap between the trial population and this more meaningful target population. Our procedure allows the analyst to improve overlap by modifying the target population for only a subset of covariates, rather than modifying all covariates or discarding observations, which is inconsistent with the goal of estimating the effect in a clinically relevant population. To adapt Problem 5 for transportability, we first introduce some additional notation. Let $R_{i}=r$ be an indicator of whether an observation belongs to the trial population ($r=1$) or the target population ($r=0$). We only observe $Y_{i}$ and $Z_{i}$ when $R_{i}=1$ but observe $\bm{X}_{i}$ for both populations. Let $n_{r}$ indicate the sample size in population $r$. Given additional assumptions outlined in Dahabreh et al. (2020), the transported effect estimand is identifiable with observed data. We propose the following modification of Problem (5) for transportability,
@@
Then, our proposed estimator for the transported effect is $\tau_{w^{*t}}=\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}Z_{i}R_{i}Y_{i}-\frac{1}{n}\sum_{i=1}^{n}w^{*t}_{i}(1-Z_{i})R_{i}Y_{i}$. The interpretation of $c(\bm{x})$ and $g(\bm{x})$ remains the same as described in Section 3.2, however the selection of $g(\bm{x})$ may be of even more importance in order to maintain a target population of interest.
Asymptotic properties of the proposed estimator
We derive asymptotic results for the solution to Equation (6) which we then use to develop asymptotic results for our proposed estimator, $\hat{\tau}_{w^{*}}$, under various conditions. All proofs of results in this section are in supplementary Section 1. We first require additional regularity conditions, similar to those in Hirano et al. (2003); Chan et al. (2016) and Källberg and Waernbaum (2023), on the set balancing functions, $b(\bm{x})$.
Assumption 3: Balance function conditions.
For the remainder of this section, we notate $w_{i}^{*}$ as $w^{*}(\bm{X}_{i})$ to explicitly show the relationship to $\bm{X}_{i}$. Given Assumptions Assumption 1: Strong ignorability, Assumption 2: Relaxed positivity, and Assumption 3: Balance function conditions, $w_{1}^{*}(\bm{X}_{i})$ and $w_{0}^{*}(\bm{X}_{i})$ are consistent for their asymptotic counterparts, $\tilde{w}_{1}(\bm{X}_{i})$ and $\tilde{w}_{0}(\bm{X}_{i})$, which uniquely solve the population version of Equation (6) (supplementary Section 1, Proposition 5); a majority of the asymptotic results follow from this fact. Since Problem (5) implies a single propensity score model, our proposed estimator is doubly robust as stated below in Theorem 3.
Theorem 3.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 3: Balance function conditions hold and $\rho^{\prime\prime}\{\cdot\}$ is continuous over $\mathbb{R}$. Assume that the expectation of $c(\bm{X})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then, $\hat{\tau}_{w^{*}}$ is doubly robust in the sense that if either 1) Assumption Assumption 2: Positivity holds and $e(\bm{x})$ satisfies conditions (7a)-(7c) or; 2) Assumption Assumption 2: Relaxed positivity holds, $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$, then $\hat{\tau}_{w^{*}}$ is consistent for the ATE.
Note that $\rho^{\prime\prime}\{\cdot\}$ is continuous for common dispersion measures such as $w^{2}$ and $w\log(w)$. In contrast, $\hat{\tau}_{w^{bw}}$ is not doubly robust (Chan et al., 2016). Since our estimator is motivated by scenarios where positivity does not hold, the second condition that guarantees consistency (i.e., $\tau(\bm{x})$ is linear in only $c(\bm{x})$) is of primary importance. Let $\tau_{g,\tilde{w}}(\bm{x},z)=(\bm{\beta}_{1}-\bm{\beta}_{0})^{T}c(\bm{x})+\bm{\lambda}_{1}^{T}\tilde{w}_{1}(\bm{x})zg(\bm{x})-\bm{\lambda}_{0}^{T}\tilde{w}_{0}(\bm{x})(1-z)g(\bm{x})$ for some $\bm{\beta}_{z}\in\mathbb{R}^{K}$ and $\bm{\lambda}_{z}\in\mathbb{R}^{L}$. When $\tau(\bm{x})$ is linear in both $c(\bm{x})$ and $g(\bm{x})$, our proposed estimator is consistent for the modified estimand $\tau_{g,\tilde{w}}=E[E[Y|Z=1,c(\bm{X}),\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]]-E[E[Y|Z=0,c(\bm{X}),\tilde{w}_{0}(\bm{X})\{1-e(\bm{X})\}g(\bm{X})]]=E[\tau_{g,\tilde{w}}(\bm{X},Z)]$ as stated in Corollary 3.1. Note that $\tau_{g,\tilde{w}}$ is indeed a causal estimand because $E[\tilde{w}_{1}(\bm{X})e(\bm{X})g(\bm{X})]=E[\tilde{w}_{0}(\bm{X})\{1-e(\bm{X})\}g(\bm{X})]$ is a first order condition of the population version of Equation (6).
Corollary 3.1.
Suppose Assumptions Assumption 1: Strong ignorability and Assumption 2: Relaxed positivity hold and $\rho^{\prime\prime}\{\cdot\}$ is continuous over $\mathbb{R}$. Assume that the expectation of $c(\bm{X})$ exists and that $\text{Var}\{Y(a)\}<\infty$ for $a=0,1$. Then, $\hat{\tau}_{w^{*}}$ is consistent for $\tau_{g,\tilde{w}}$ if $e(\bm{x})$ does not satisfy conditions (7a)-(7c) and $\mu_{z}(\bm{x})$ and $\tau(\bm{x})$ are both linear in $c(\bm{x})$ and $g(\bm{x})$.
When positivity does hold and all models implied by Problem 5 are correctly specified, our proposed estimator is semiparametrically efficient as stated in Theorem 4.
Theorem 4.
Suppose Assumptions Assumption 1: Strong ignorability, Assumption 2: Positivity, and Assumption 3: Balance function conditions hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ satisfies conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then,
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt})$
where $V_{opt}=E\left[\frac{\text{Var}(Y(1)|\bm{X})}{e(\bm{X})}+\frac{\text{Var}(Y(0)|\bm{X})}{1-e(\bm{X})}+\{\tau(\bm{X})-\tau\}^{2}\right]$ is the semiparametric efficiency bound (Hahn, 1998).
However, it is rare that the implied propensity score model is correctly specified and our proposed estimator is motivated by scenarios where Assumption Assumption 2: Positivity does not hold. Thus, we derive and present the asymptotic variance of our estimator when only the outcome and models are correctly specified in Corollary 4.1.
Corollary 4.1.
Suppose Assumptions Assumption 1: Strong ignorability, Assumption 2: Relaxed positivity, and Assumption 3: Balance function conditions hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Suppose $e(\bm{x})$ does not satisfy conditions (7a)-(7c), $\mu_{z}(\bm{x})$ is linear in $c(\bm{x})$ and $g(\bm{x})$, and $\tau(\bm{x})$ is linear in $c(\bm{x})$. Then,
$\sqrt{n}(\hat{\tau}_{w^{*}}-\tau)\to N(0,V_{opt}^{*})$
where $V_{opt}^{*}=E[\tilde{w}_{1}(\bm{X})^{2}e(\bm{X})\text{Var}\{Y(1)|\bm{X}\}+\tilde{w}_{0}(\bm{X})^{2}\{1-e(\bm{X})\}\text{Var}\{Y(0)|\bm{X}\}+\{\tau(\bm{X})-\tau\}^{2}]$.
Of key importance, this variance result only requires the relaxed positivity assumption, rather than the standard positivity assumption. However, to compare $V_{opt}$ and $V^{*}_{opt}$, consider the case where positivity does hold. Since the asymptotic weights tend to be less extreme (i.e., $\tilde{w}_{1}(\bm{X}_{i})\leq\{e(\bm{X}_{i})\}^{-1}$ and $\tilde{w}_{0}(\bm{X}_{i})\leq\{1-e(\bm{X}_{i})\}^{-1}$) than IPW through the restriction on the implied propensity score model, $V^{*}_{opt}\leq V_{opt}$. Thus, our proposed estimator does indeed have reduced asymptotic variance, especially in scenarios where $e(\bm{x})$ is close to zero or positivity does not hold. Specifically, as the $g(\bm{x})$ set gets larger, $\tilde{w}_{z}(\bm{x})$ become less extreme and correspondingly $V^{*}_{opt}$ becomes smaller. We further derive the asymptotic variance of our proposed estimator with respect to the modified estimand, $\tau_{g,\tilde{w}}$, when $\tau(\bm{x})$ is a linear function of both $c(\bm{x})$ and $g(\bm{x})$ in Corollary 4.2.
Corollary 4.2.
Suppose Assumptions Assumption 1: Strong ignorability, Assumption 2: Relaxed positivity, and Assumption 3: Balance function conditions hold and $\text{Var}\{Y(a)|\bm{X}=\bm{x}\}<\infty$ for all $\bm{x}$ and $a=0,1$. Assume that $\rho^{\prime\prime}\{\cdot\}$ and $\rho^{\prime\prime\prime}\{\cdot\}$ are continuous over $\mathbb{R}$. Suppose $e(\bm{x})$ does not satisfy conditions (7a)-(7c) and $\mu_{z}(\bm{x})$ and $\tau(\bm{x})$ are linear in both $c(\bm{x})$ and $g(\bm{x})$. Then,
[…]
where $V_{g,\tilde{w}}=E[\tilde{w}_{1}(\bm{X})^{2}e(\bm{X})\text{Var}\{Y(1)|\bm{X}\}+\tilde{w}_{0}(\bm{X})^{2}\{1-e(\bm{X})\}\text{Var}\{Y(0)|\bm{X}\}+\{\tau_{g,\tilde{w}}(\bm{X},Z)-\tau_{g,\tilde{w}}+r_{1}(\bm{X},Z)-r_{0}(\bm{X},Z)\}^{2}].$
The functions $r_{z}(\bm{x},z)$ account for the difference between the estimated and asymptotic weights. To avoid introducing substantial additional notation, we present the exact forms of these functions in the proof of Corollary 4.2 in supplementary Section 1. Note that the additional assumptions on $\rho\{\cdot\}$ hold for common dispersion measures such as $w^{2}$ and $w\log(w)$. To compare to Corollary 4.1, consider the case where the assumptions of Corollary 4.2 hold and the direct balancing weights estimator is computed. In this case, the asymptotic variance of the direct balancing estimator takes the same form as $V_{opt}^{*}$ but with the asymptotic counterparts of direct balancing weights; we will reference this variance as $V_{opt}^{bw}$. Our proposed asymptotic weights are less extreme than the asymptotic direct balancing weights, however, $\text{Var}\{\tau_{g,\tilde{w}}(\bm{X},Z)\}\gtrsim\text{Var}\{\tau(\bm{X})\}$ as $\text{Var}\{w_{z}(\bm{X})I[Z=z]g(\bm{X})\}\gtrsim\text{Var}\{g(\bm{X})\}$. Thus, there is an unclear relationship between $V_{g,\tilde{w}}$ and $V^{bw}_{opt}$; in some scenarios $V_{g,\tilde{w}}>V^{bw}_{opt}$ while in others $V_{g,\tilde{w}}<V^{bw}_{opt}$. To further explore this relationship, we perform a simulation study to compare the asymptotic variance of these estimators (supplementary Section 4). When overlap is poor and there is low to moderate treatment effect heterogeneity, $V_{g,\tilde{w}}$ tends to be smaller than $V^{bw}_{opt}$ (supplementary Figure 1). With the addition of the constraints $w_{i}\geq 0$ and $\sum_{i=1}^{n}w_{i}Z_{i}=\sum_{i=1}^{n}w_{i}(1-Z_{i})$ to both optimization problems, the direct balancing weights are even more extreme such that $\text{Var}(\hat{\tau}_{w^{*}})\lesssim\text{Var}(\hat{\tau}_{w^{wb}})$ across a majority of data generating scenarios (supplementary Figure 3); in fact, when overlap is poor $\text{Var}(\hat{\tau}_{w^{*}})<<\text{Var}(\hat{\tau}_{w^{wb}})$.
@@
To select the optimal $g(\bm{x})$ set, in terms of error with respect to the ATE, one approach is to examine all possible covariate function sets and select the set that 1) yields a solution to Problem (5); and 2) minimizes term (8a). However, examining all possible covariate subsets is computationally prohibitive, especially as covariate dimension increases. Therefore, we propose a greedy algorithm (Algorithm 1) for identifying an approximately optimal $g(\bm{x})$ subset according to a specific metric aligned with either the model- or design-based perspective on estimator error. We also propose an additional algorithm (supplementary Algorithm 1) that is less computationally intensive than Algorithm 1; we found these algorithms resulted in estimators with similar performance when there was no correlation between covariates (supplementary Figure 4). However, we generally recommend the use of Algorithm 1 if it is not computationally prohibitive.
@@
1:Inputs: The matrix $\mathbf{X}_{n\times p}$, set of covariate functions $b_{1}(\bm{X}),\ldots,b_{J}(\bm{X})$, and a metric $m$ that can be computed for each $b_{j}(\bm{X})$
@@
3:while $\{w^{*}_{i}\}_{i=1}^{n}$ given $g(\bm{X})=b(\bm{X})^{-}$ does not exist do
4:  Compute $m$ for $\{b_{1}(\bm{X}),\ldots,b_{j}(\bm{X})\}\setminus b(\bm{X})^{-}\}$ and identify $b^{*}(\bm{X})$, the covariate with the minimum or maximum value of the metric $m$
@@
6:  Solve Problem (5) with $g(\bm{X})=b(\bm{X})^{-}$ and $c(\bm{X})=\{b_{1}(\bm{X}),\ldots,b_{j}\}\setminus b(\bm{X})^{-}$